### v3

# L-MARS: Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search

Abstract
We present L-MARS (Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search), a multi-agent retrieval framework for grounded legal question answering that decomposes queries into structured sub-problems, retrieves evidence via agentic web search, filters results through a verification agent, and synthesizes cited answers. Existing legal QA benchmarks test either closed-book reasoning or retrieval over fixed corpora, but neither captures scenarios requiring current legal information. We introduce LegalSearchQA, a 50-question benchmark across five legal domains whose answers depend on recent developments that post-date model training data. L-MARS achieves 96.0% accuracy in LegalSearchQA, a 38.0% improvement over zero-shot (58.0%), while chain-of-thought prompting degrades performance to 30.0%. On Bar Exam QA [28], a reasoning-focused benchmark of 594 bar examination questions, retrieval provides negligible gains (+0.7 percentage points), consistent with prior findings. These results show that agentic retrieval dramatically improves legal QA when tasks require up-to-date factual knowledge, but the benefit is benchmark-dependent, underscoring the need for retrieval-focused evaluation. Code and data: https://github.com/boqiny/L-MARS.
1 Introduction
Large language models (LLMs) such as GPT [1], Claude [3], and Gemini [4] are increasingly applied to legal tasks including case law retrieval [26], statutory interpretation, and automated legal assistance [2]. However, direct application often produces hallucinations [5, 27] where the model states confidently incorrect legal facts. In legal contexts such errors carry real-world risk, as wrong citations or outdated statutes can undermine credibility and affect decision-making.
Domain-specific fine-tuning [6] can help but requires frequent retraining as laws change, making it impractical for continuously evolving domains. Retrieval-Augmented Generation (RAG) [7] avoids retraining by fetching relevant documents at inference time, but its effectiveness depends on retrieval quality. Recent work has shown that on reasoning-intensive legal benchmarks, standard retrieval pipelines barely improve downstream QA accuracy. On Bar Exam QA, even the best traditional retriever improves GPT-4o-mini accuracy by only 0.5 percentage points over zero-shot [28].
We introduce L-MARS, a multi-agent retrieval framework that addresses these challenges through structured agent orchestration. L-MARS combines online search via the Serper API [8], offline retrieval using BM25 over local documents, and case law retrieval through the CourtListener API [9]. Unlike naive retrieve-and-generate pipelines, L-MARS employs a multi-agent workflow that decomposes queries into structured search intents, executes targeted retrieval, filters evidence through a Judge Agent, and synthesizes grounded answers.
Figure 1: Comparison of the two L-MARS operating modes. (a) Simple Mode: executes a single-pass retrieval–summarization pipeline. (b) Multi-Turn Mode: adds Judge Agent–guided iterations with sufficiency checks and query refinement.
L-MARS operates in two modes. In Simple Mode, a single-pass pipeline decomposes the query, retrieves and filters evidence, and synthesizes a grounded answer. In Multi-Turn Mode, an iterative loop enables the system to adaptively search, verify, and refine until an evidence sufficiency threshold is met. In this work we evaluate Simple Mode to isolate the contribution of agent orchestration from iterative refinement.
In legal practice, outdated information can lead to materially incorrect advice. Our results suggest that retrieval is essential for legal QA tasks that depend on current facts, and may be important for reliable deployment in domains where knowledge changes rapidly.
Our contributions are threefold. First, we introduce LegalSearchQA, a benchmark for legal QA requiring post-training, time-sensitive knowledge. Second, we present L-MARS, a multi-agent retrieval framework for grounded legal question answering. Third, we show empirically that retrieval yields large gains only in retrieval-dependent settings, while chain-of-thought reasoning can degrade performance when parametric knowledge is outdated. All prompts, evaluation scripts, and raw outputs are released to ensure full reproducibility.
2 Related Work
Agentic RAG and Search Systems
RAG grounds model output in external evidence to reduce hallucination [7], and interleaving reasoning with retrieval [30, 10] further improves factual grounding. Multi-agent frameworks [11, 12] enable role specialization and iterative critique, with recent agentic RAG variants combining planning, verification, and retrieval for long-horizon tasks [31, 33]. On the search side, Search-o1 and WebThinker [13, 14] integrate agentic search into reasoning models, while OpenAI’s Deep Research [15] exposes a production-grade web-scale research loop. These systems complement reasoning models [16, 17] that scale deliberation at test time through chain-of-thought prompting [18], self-consistency [19], verifier-based calibration [32], and dedicated reasoning architectures [20, 21].
Legal AI and Benchmarks
Legal AI evaluation spans benchmarks for reasoning [24], multi-task NLU [22], Chinese legal knowledge [34], and pretraining data [23]. More recent work targets retrieval and reasoning specifically: Bar Exam QA [28] captures realistic legal RAG settings where standard retrieval fails, and LEXam [29] benchmarks multi-step legal reasoning across languages. On the systems side, ChatLaw [35] explores multi-agent approaches for legal tasks. We note that many LegalBench [24] subtasks are now saturated, with GPT-4o-mini achieving 100% on the rule_qa, limiting their discriminative value for knowledge-intensive legal QA.
3 Methodology
3.1 Problem Formation
We formalize legal question answering as generating both a reasoning chain $R$ and a final answer $a$ given a user query $q$, external retrieval results $D$, and optional user clarifications $U$. Each query state is represented as $(q,U,D)$, which evolves iteratively through the workflow. At each step, the system decides whether (i) expand $D$ by issuing new search queries, or (ii) terminate with a sufficient answer. Formally, the system computes a mapping
$(q,U,D)\;\mapsto\;(R,a),$
where $R=\{r_{1},r_{2},\dots,r_{T}\}$ is a sequence of reasoning steps interleaved with the retrieved evidence, and $a$ is the final response string. The Judge Agent makes a sufficiency decision $s\in\{\textsc{Sufficient},\textsc{Insufficient}\}$ and refinement notes $E$ guiding the next query state. The loop terminates when $s=\textsc{Sufficient}$ or after $M$ iterations, as illustrated in Figure 1.
3.2 Workflow and Agents Overview
Workflow
L-MARS is a multi-agent workflow implemented in LangGraph with structured output validation and tool use. The system employs a directed graph in which agents are represented as nodes and the control flow is represented as edges with conditional routing. The workflow maintains a centralized WorkflowState that tracks: (i) the evolving query representation, (ii) the accumulated search results with metadata, (iii) the iteration history and refinement notes, and (iv) the intermediate agent outputs. State transitions are deterministic and type-checked, ensuring reproducibility across runs.
Agents
The Query agent parses the user’s question into a structured query result. The Search Agent executes tool calls to the retrieval backends and returns normalized SearchResult objects (title, url, snippet, optional full content, source type). The Judge Agent performs evidence sufficiency checks and an explicit checklist (factual support, jurisdiction match, temporal specificity, and contradiction analysis). Summary Agent composes the final response with citations and rationale. Detailed agent implementations can be found in the Appendix A.
3.3 Agentic Search
Online Search
We adopt an agentic retrieval policy inspired by Search-o1[13], interleaving targeted web queries with stepwise reasoning rather than performing a single problem-oriented fetch. Concretely, the Search Agent triggers retrieval when either (i) the Judge Agent flags a knowledge gap (e.g. missing statute/case authority, date or jurisdiction ambiguity), or (ii) the query classifier predicts that external evidence will reduce answer uncertainty. Once activated, the agent executes one of two complementary search modes against Google Serper, as shown in Table 1 :
Basic Search (titles/snippets) parses organic results into a normalized schema (title, URL, site, date, snippet). This mode is fast and used for query exploration, de-duplication, and document selection under a strict token budget.
Enhanced Search (content extraction) uses the top-$m$ URLs selected by the agent (default m=3m{=}3), then fetches page contents and PDFs with robust headers and timeouts. HTML is cleaned with BeautifulSoup; PDFs are parsed with pdfplumber. To avoid flooding the context, we perform a snippet-anchored extraction: given the search snippet, we locate the best-matching sentence in the full text using a token-level $F_{1}$ overlap and return a fixed window of the surrounding context (2.5k chars). This preserves local evidence while discarding boilerplate and navigation chrome. The extractor also normalizes whitespace, strips punctuation, and truncates long pages to a hard cap to guarantee bounded latency.
Table 1: Comparison of Basic vs. Enhanced (Deep) Search on the query “Can I work remotely in the United States as an F1 student as of 2025?”
Dimension Basic Search (Titles & Snippets) Enhanced Search (Full Content)
Source Types Interstride blog, Reddit, Quora Interstride blog (full text), others truncated
Output Length 1,145 chars, 19 lines 4,154 chars, 15 lines
Example Result “Even if you are working remotely… must have work authorization.” “Remote work for international students on F-1 visas: Any work done on US soil requires authorization…”
Content Depth Snippets only; limited context Full article text (4k cap); snippet-anchored window (2.5k chars)
Latency Fast Slower; scraping + parsing
Use Case Quick exploration, de-duplication Detailed evidence; grounding in authoritative text
Local RAG Indexing
We also maintain a local retrieval index over user-provided documents. Markdown files are segmented into overlapping windows of 500 characters with 100-character stride and indexed using the BM25 ranking function [25]. At query time, the retriever returns the top-$k$ segments (default $k=5$), enriched with document titles and section headers to preserve context. The indexing process is dynamic: new files added to the inputs directory are automatically detected and incorporated without requiring a system restart.
CourtListener Integration
To incorporate authoritative case law into the reasoning process, we integrate the CourtListener Legal Search API as a tool, providing access to millions of judicial opinions, oral arguments, and dockets curated by the Free Law Project. When enabled, the Search Agent queries by party name, citation, docket number, or keyword. The API returns structured metadata (case name, citation, court, filing date, opinion text), which we normalize into a uniform schema. This integration allows the system to ground answers in both retrieved secondary sources and primary case law to improve legal credibility.
3.4 L-MARS Operating Mode
Simple Mode
The Simple Mode executes a single-pass pipeline that prioritizes minimal latency. The Query Agent generates structured intents from the user question, which the Search Agent uses to retrieve evidence from enabled sources (Serper web snippets, optional BM25 index, and CourtListener). The Summary Agent then composes an answer with citations in one step.
Multi-Turn Mode
The Multi-Turn mode adds an iterative search–judge–refine loop to improve recall and reduce hallucinations. The Query Agent first proposes clarifying follow-ups; optional user responses are folded into a structured query state. Each iteration generates targeted search queries, runs deep search (top-$3$ with content), and invokes a deterministic Judge Agent (T=0T{=}0) that performs chain-of-thought analysis, source quality checks, and date/jurisdiction/contradiction screening to decide sufficiency. If insufficient, the system refines the query state and repeats up to a fixed maximum. A final Summary Agent composes the answer.
Judge Agent and Stopping Rule
The Judge Agent runs with temperature =0=0 to reduce variance and improve reproducibility. Its prompt elicits three outputs: (i) a chain-of-thought style rationale, (ii) a structured checklist decision on sufficiency, and (iii) missing-evidence directives that guide the next refinement queries. The loop ends once the decision is made Sufficient or when the maximum iteration limit is reached. In practice, we found the Judge Agent especially useful when search results included unofficial forums (e.g., Reddit, Quora) instead of government domains; in such cases it flagged insufficiency and steered subsequent searches toward more compelling, authoritative sources such as government websites and official regulations.
We describe both operating modes and all three retrieval backends for completeness, as they form the full L-MARS system available in our codebase. In our evaluation we use only Simple Mode with Serper web search, since this minimal configuration already achieves 96% accuracy on LegalSearchQA. Local RAG indexing and CourtListener case law retrieval are available as optional backends for practitioners working with proprietary document collections or case-specific research, but are not exercised in our benchmark evaluation.
4 Experiments
4.1 Tasks and Datasets
LegalSearchQA
Existing legal QA benchmarks test either closed-book reasoning on static exam questions or retrieval over fixed corpora. Neither reflects the real-world setting where a legal professional needs current information, such as recently enacted statutes, updated regulatory thresholds, or new court decisions, that was not available when the model was trained. To address this gap we introduce LegalSearchQA, a 50-question multiple-choice benchmark designed to evaluate retrieval-augmented legal QA systems on current legal knowledge.
Each question has four answer choices and exactly one correct answer. Questions span five legal domains: Tax, Corporate & Financial Regulation (13), Labor & Employment (11), Immigration (9), Technology & Privacy (9), and Criminal, Drug & State Law (8). The benchmark targets three categories of legal knowledge: status questions about the current state of laws, threshold questions about specific numerical values such as dollar amounts or deadlines, and requirement questions about what rules mandate. All questions were authored by the paper’s authors and verified against primary legal authorities such as IRS.gov, USCIS.gov, SEC.gov, and Supreme Court opinions during March 2026. By construction, correct answers require information from 2024 to 2026 that is unlikely to appear in model training data, making web retrieval essential. To reduce author bias, each question is grounded in a verifiable primary legal source and the task is purely factual retrieval rather than subjective interpretation, limiting degrees of freedom in answer labeling. While the absolute accuracy of 96% is high, this reflects the benchmark’s design: questions are intentionally selected to isolate retrieval-dependent knowledge where access to current information provides a clear advantage.
To further reduce potential bias, we release all questions with source citations, timestamps, and verification notes. Each question is constructed to have a single unambiguous factual answer grounded in primary legal sources (e.g., IRS, SEC, Federal Register). Because the task is factual rather than interpretive, correctness can be independently verified without subjective judgment. The complete dataset is available at https://huggingface.co/datasets/boqiny/LegalSearchQA.
Bar Exam QA
We additionally evaluate on Bar Exam QA [28], a legal QA benchmark drawn from the Reasoning-Focused Legal Retrieval Benchmark. The dataset consists of multistate bar examination questions that test legal reasoning across core areas of U.S. law. We use the publicly released Historical MBE subset of 594 questions with subject labels spanning six domains: Contracts (113), Torts (112), Constitutional Law (95), Evidence (93), Real Property (92), and Criminal Law (89).
The authors of Bar Exam QA showed that standard retrieval pipelines achieve very low recall on this benchmark and their best retriever improves GPT-4o-mini accuracy by only 0.5 percentage points over zero-shot. We include this benchmark to confirm that retrieval provides negligible gains on reasoning-focused tasks. We note that the Historical MBE subset overlaps with the MMLU auxiliary training set [36] and may inflate zero-shot baselines through memorization; we mitigate this by focusing primary claims on LegalSearchQA, where memorization is not a factor.
4.2 Baselines and Experimental Conditions
We compare three conditions, all using GPT-4o-mini as the backbone model:
Zero-Shot
The model answers directly with no retrieval and no explicit reasoning instructions. The prompt presents the question and four choices, instructing the model to select a single answer letter.
Chain-of-Thought (CoT)
The model is prompted to reason step-by-step: identify the legal issue, recall the relevant rule, apply it to the facts, and eliminate incorrect options before selecting an answer. No external retrieval is provided.
L-MARS (Simple Mode)
The full L-MARS pipeline in single-pass mode: the Query Agent decomposes the question into structured search intents, the Search Agent retrieves from Serper with enhanced content extraction, and the Summary Agent synthesizes the final answer with citations. All agents use GPT-4o-mini with T=0T{=}0.
This experimental ladder isolates specific contributions: Zero-Shot to CoT tests whether reasoning alone helps, and CoT to L-MARS tests whether agentic retrieval adds value beyond reasoning. We leave evaluation of the Multi-Turn Mode to future work, as it targets open-ended legal research questions where a single retrieval pass may not surface sufficient evidence, a setting not captured by our current multiple-choice benchmarks.
While L-MARS includes multiple components, our evaluation isolates the contribution of retrieval itself by comparing zero-shot against the full pipeline. Since Simple Mode excludes iterative refinement and Judge Agent loops, the observed improvements can be attributed primarily to structured web retrieval rather than multi-turn reasoning.
4.3 Implementation Details
All conditions were evaluated using a standardized pipeline. For LegalSearchQA, we evaluated all 50 questions under each of the three conditions. For Bar Exam QA, we evaluated the full 594-question subset under zero-shot and L-MARS conditions. For the zero-shot and CoT conditions, prompts instruct the model to select a single answer letter (A–D). For L-MARS, all agents use GPT-4o-mini with T=0T{=}0; CourtListener integration was disabled since neither benchmark requires specific case citations. On Bar Exam QA, we additionally tested L-MARS with the Tavily search backend as an alternative to Serper. The complete evaluation pipeline, including prompts, scripts, and all results, is available at https://github.com/boqiny/L-MARS.
5 Results
5.1 LegalSearchQA
Table 2 reports accuracy across three conditions on LegalSearchQA. L-MARS achieves 96.0% accuracy, improving over zero-shot by 38 percentage points while maintaining moderate latency at 2.4s per question. Chain-of-thought prompting degrades performance from 58.0% to 30.0%, falling below the 25% random baseline for four-choice questions. Without access to current information, CoT causes the model to construct plausible but outdated rationales and arrive at wrong answers with high confidence. We term this confident confabulation: the model recalls stale figures from prior tax years or superseded regulations, then builds step-by-step justifications around them. L-MARS rescued 19 questions that zero-shot answered incorrectly while regressing on none. We use a standard legal reasoning CoT prompt following the issue-rule-application-conclusion (IRAC) structure. We attribute the degradation to reliance on outdated parametric knowledge rather than prompt design, as the model produces internally coherent reasoning from stale premises across all five domains. We verified consistent degradation patterns with temperature T=0T{=}0 across all five domains, suggesting the effect is robust rather than an artifact of a single prompt configuration.
Table 2: Accuracy on LegalSearchQA (50 questions). All conditions use GPT-4o-mini.
Condition Retrieval Accuracy (%) Correct/Total Avg. Latency (s)
Zero-Shot None 58.0 29/50 0.7
Chain-of-Thought None 30.0 15/50 6.8
L-MARS (Simple) Serper + agents 96.0 48/50 2.4
Figure 2 and Table 3 break down accuracy by legal domain. The largest gains appear in Tax, Corporate & Financial Regulation, where L-MARS improves over zero-shot by 61.5 percentage points. Questions in this domain ask for specific dollar amounts, filing deadlines, and regulatory thresholds that changed in 2025 and 2026. Zero-shot accuracy here is only 30.8%, reflecting the model’s inability to access post-training information; L-MARS reaches 92.3% by retrieving current IRS and SEC sources. The smallest gains appear in Criminal, Drug & State Law, where zero-shot already achieves 87.5% because many questions concern well-publicized events that likely entered the model’s training data through news coverage. Notably, CoT falls below the random baseline in three of five domains, confirming that reasoning from stale knowledge is worse than no reasoning at all when current facts are required.
Figure 2: Per-domain accuracy on LegalSearchQA. L-MARS achieves near-perfect accuracy across all domains. CoT falls below the 25% random baseline in three of five domains.
Table 3: Per-domain accuracy (%) on LegalSearchQA. $\Delta$ is L-MARS gain over zero-shot.
Domain N Zero-Shot CoT L-MARS $\Delta$
Tax, Corporate & Financial 13 30.8 15.4 92.3 +61.5
Labor & Employment 11 63.6 18.2 100.0 +36.4
Technology & Privacy 9 55.6 44.4 100.0 +44.4
Immigration 9 66.7 22.2 88.9 +22.2
Criminal, Drug & State Law 8 87.5 62.5 100.0 +12.5
Overall 50 58.0 30.0 96.0 +38.0
5.2 Bar Exam QA
Figure 3 visualizes the contrast between the two benchmarks. On Bar Exam QA (Table 4), L-MARS provides a negligible 0.7 point gain with Serper and a slight 0.8 point degradation with Tavily. Bar Exam QA tests legal reasoning over hypothetical fact patterns whose answers follow from general legal principles already in the model’s parameters, so web retrieval adds noise rather than signal. The gap between a 38 point gain on LegalSearchQA and a 0.7 point gain on Bar Exam QA confirms that the value of agentic retrieval depends on whether the task requires current information or abstract reasoning.
Figure 3: Benchmark comparison. On retrieval-dependent LegalSearchQA, L-MARS gains 38 percentage points over zero-shot. On reasoning-focused Bar Exam QA, retrieval provides negligible improvement.
Table 4: Accuracy on Bar Exam QA (594 questions). All conditions use GPT-4o-mini.
Condition Acc. (%) Correct
Zero-Shot 55.2 328/594
L-MARS (Serper) 55.9 332/594
L-MARS (Tavily) 54.4 323/594
Prior work [28]:
GPT-4o-mini 49.7
+ Best retriever 50.2
Table 5: Per-subject accuracy (%) on Bar Exam QA.
Subject N ZS L-MARS
Const. Law 95 81.1 76.8
Contracts 113 46.9 52.2
Criminal 89 43.8 41.6
Evidence 93 54.8 61.3
Real Prop. 92 47.8 50.0
Torts 112 57.1 53.6
Table 5 shows per-subject results. L-MARS with Serper yields modest gains on Contracts (+5.3 points) and Evidence (+6.5 points) but slight regressions on Constitutional Law and Torts. No subject shows a large or consistent benefit from retrieval, reinforcing the conclusion that this benchmark primarily tests reasoning ability rather than information access.
5.3 Error Analysis
L-MARS answered only two LegalSearchQA questions incorrectly. Q27 asked about H-1B weighted selection details involving a nuanced procedural rule about wage-level assignment for multiple registrations; the retrieved results did not surface the specific Federal Register provision. Q44 asked for the exact effective date of T+1 settlement; search results discussed the policy broadly but missed the precise date. Both errors share a pattern: the correct answer requires a specific fact buried in a longer document, and snippet-level retrieval missed the critical sentence.
The CoT degradation deserves separate attention. On tax threshold questions, for instance, the model recalls outdated dollar amounts from prior years and builds multi-step justifications around them. This pattern is not unique to legal QA. In any domain where facts change faster than model training cycles, prompting the model to reason more carefully can amplify rather than correct errors, because the model reasons from wrong premises with greater conviction. This finding suggests caution when deploying CoT in fast-changing domains without concurrent retrieval.
5.4 Ablation: Retrieval Depth
[…]
6 Conclusion
We presented L-MARS, a multi-agent retrieval framework for legal QA, and LegalSearchQA, a 50-question benchmark requiring current legal knowledge that post-dates model training data. L-MARS achieves 96.0% on LegalSearchQA, a 38 percentage point gain over zero-shot, while CoT degrades accuracy to 30.0% through confident confabulation from stale knowledge. On reasoning-focused Bar Exam QA, retrieval provides negligible gains, confirming that the value of agentic retrieval is benchmark-dependent. These findings underscore the need for retrieval-focused legal benchmarks and suggest caution when deploying CoT in fast-changing domains without concurrent retrieval.
Limitations
LegalSearchQA contains 50 questions verified as of March 2026; the benchmark is sufficient for demonstrating large effect sizes but limits per-domain statistical power, and answers may become outdated as laws change. We evaluate only Simple Mode with web search; the Multi-Turn Mode, local RAG indexing, and CourtListener integration are available in the released codebase but not benchmarked here, as the simple configuration already demonstrates the core finding. Our evaluation covers U.S. law in English with a single backbone model, and performance remains bounded by search engine quality, as seen in the two errors where snippet-level retrieval missed critical details buried in longer documents.
Appendix
Appendix A Agent Instructions
Query Agent
The Query Agent is a stateless pass-through that wraps the user
query into a structured QueryGeneration dataclass:
    @dataclass
    class QueryGeneration:
        query: str
        query_type: str = "web_search"
        priority: str = "high"
    class QueryAgent:
        def build_query(self, user_query: str) -> QueryGeneration:
            return QueryGeneration(query=user_query.strip())
Summary Agent
You are a bar exam expert. A user will provide retrieved legal
sources and a multiple-choice question. Use the retrieved sources
together with your legal knowledge to select the single best answer.
Rules:
1. Your response MUST start with exactly one letter on the first
   line: A, B, C, or D.
2. Follow with a concise 1-2 sentence rationale citing the
   controlling legal rule.
3. Bar exam questions test specific exceptions and nuances -- do
   not just apply the most common general rule if the specific
   facts call for an exception.
Response format:
[Letter]
[Brief rationale]
The user message is structured as:
  Retrieved Evidence:
  <concatenated search results with source, title, content>
  Question:
  <full MCQ with answer choices A-D>
Judge Agent System Prompt
You are a legal research quality judge for US bar exam questions.
You will receive a bar exam multiple-choice question and a list of
retrieved search result snippets. Your job is to decide whether the
retrieved evidence is sufficient to reliably answer the question.
"Sufficient" means: at least one result contains specific legal
rules, statutes, case holdings, or doctrinal exceptions that
directly address the core legal issue tested by the question -- not
just a restatement of the facts.
If the evidence is NOT sufficient, provide a better search query
that targets primary legal sources (e.g., specific rule names,
statute citations, or doctrine names) rather than generic question
text.
Respond with valid JSON only, no markdown fences:
{"sufficient": true_or_false, "reason": "one sentence",
 "next_query": "refined query (omit if sufficient)"}
Pre-filtering: Before LLM evaluation, results matching any of
these patterns are deterministically removed at zero cost:
  HTTP error, 403/404, Connection error, Request timed out,
  JavaScript is disabled, No search results found, etc.
Appendix B Case Study: Post-Training Knowledge Gap
We present two representative examples from LegalSearchQA where zero-shot and chain-of-thought both fail due to outdated parametric knowledge, while L-MARS retrieves the correct current information.
Q1 (lsqa_001): Executive Order 14110 Status
Question: What is the current status of Executive Order 14110, the
Biden administration’s executive order on AI safety signed in
October 2023?
A. It remains in full effect and is being actively enforced
B. It was revoked by President Trump on January 20, 2025
C. It was partially amended but key provisions remain
D. It expired automatically after 12 months
Correct Answer: B
Zero-Shot Response (Incorrect)
Selected: A
The model’s parametric knowledge contains EO 14110 as an active,
enforceable order -- it has no information about the January 2025
revocation, which post-dates its training data.
Chain-of-Thought Response (Incorrect)
Selected: A
The model reasons step-by-step from stale knowledge: it recalls
that EO 14110 was signed in October 2023, notes that executive
orders do not expire automatically, and concludes it must still be
in effect. The reasoning is internally consistent but factually
wrong -- an example of "confident confabulation" from outdated
parametric memory.
L-MARS Response (Correct)
Selected: B
Executive Order 14110 has been revoked, as indicated by the sources
stating that President Trump took action to revoke it during his
final week in office. Therefore, it is no longer in effect.
Retrieved sources included whitehouse.gov and Federal Register
entries confirming revocation via EO 14148 on January 20, 2025.
Q2 (lsqa_002): 2025 Standard Deduction
Question: For tax year 2025, what is the standard deduction for a
single filer under age 65, after the One Big Beautiful Bill Act
amendments?
A. \$14,600    B. \$15,000    C. \$15,350    D. \$15,750
Correct Answer: D
Zero-Shot: C (Incorrect)   CoT: C (Incorrect)   L-MARS: D (Correct)
Zero-Shot selected C (\$15,350) -- the model recalls an intermediate
figure from pre-OBBBA IRS projections but lacks knowledge of the
legislative amendment.
CoT selected C (\$15,350) -- step-by-step reasoning reinforces the
wrong figure: the model recalls the 2024 standard deduction
(\$14,600), applies an estimated inflation adjustment, and arrives
at \$15,350. The reasoning is plausible but based on stale data.
L-MARS selected D (\$15,750) -- retrieved current IRS sources
confirming the One Big Beautiful Bill Act raised the 2025 standard
deduction to \$15,750 for single filers.
B.1 Analysis
Both examples illustrate the same failure mode: the model’s parametric knowledge is confidently wrong about post-training developments. In Q1, a major policy reversal (EO revocation) is invisible to the model. In Q2, a legislative amendment changes a specific dollar figure that the model can only approximate from outdated data.
Chain-of-thought prompting amplifies these failures rather than mitigating them. By encouraging the model to reason step-by-step from stale premises, CoT produces internally coherent but factually incorrect justifications. This “confident confabulation” effect explains why CoT accuracy (30.0%) falls below even zero-shot (58.0%) on LegalSearchQA.
L-MARS resolves both cases by retrieving current authoritative sources (whitehouse.gov, IRS.gov) and grounding the answer in retrieved evidence rather than parametric memory. The system’s ability to access post-training information is the primary driver of its 38-percentage-point improvement over zero-shot on this benchmark.